# Liming Liu1    Zhenghao Xu1    Zixuan Zhang1    Hao Kang1 Zichong Li1    Chen Liang2    Weizhu Chen2    Tuo Zhao1

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable success across various domains, yet their optimization remains a significant challenge due to the complex and high-dimensional loss landscapes they inhabit. While adaptive optimizers such as AdamW are widely used, they suffer from critical limitations, including an inability to capture interdependencies between coordinates and high memory consumption. Subsequent research, exemplified by SOAP, attempts to better capture coordinate interdependence but incurs greater memory overhead, limiting scalability for massive LLMs. An alternative approach aims to reduce memory consumption through low-dimensional projection, but this leads to substantial approximation errors, resulting in less effective optimization (e.g., in terms of per-token efficiency). In this paper, we propose COSMOS, a novel hybrid optimizer that leverages the varying importance of eigensubspaces in the gradient matrix to achieve memory efficiency without compromising optimization performance. The design of COSMOS is motivated by our empirical insights and practical considerations. Specifically, COSMOS applies SOAP to the leading eigensubspace, which captures the primary optimization dynamics, and MUON to the remaining eigensubspace, which is less critical but computationally expensive to handle with SOAP. This hybrid strategy significantly reduces memory consumption while maintaining robust optimization performance, making it particularly suitable for massive LLMs. Numerical experiments on various datasets and transformer architectures are provided to demonstrate the effectiveness of COSMOS. Our code is available at https://github.com/lliu606/COSMOS.
@@ 1 Introduction
The optimization of Large Language Models (LLMs) is fundamental to their success, enabling these models to achieve state-of-the-art performance across diverse tasks. However, the high-dimensional and non-convex loss landscapes inherent to LLMs, which can contain hundreds of billions or even trillions of parameters (Brown et al., 2020; Achiam et al., 2023), present significant optimization challenges. Adaptive optimizers, such as Adam (Kingma, 2014) and its variants AdamW (Loshchilov, 2017), have emerged as de facto standards due to their ability to dynamically adjust learning rates based on the second moment of the gradient matrix. Despite their widespread adoption, these methods suffer from two critical limitations that impede their effectiveness and scalability in the context of increasingly large and complex LLMs:
  • One limitation of Adam and its variants lies in its coordinate-wie use of adaptive learning rates. While this approach dynamically adjusts learning rates for individual parameters based on their gradient history, it inherently fails to capture interdependencies between coordinates. By treating each parameter independently, this method essentially implements an efficient diagonal approximation of the preconditioning matrix. Although this reduces computational complexity, it inadequately represents the intricate curvature information of the loss landscape, particularly in the high-dimensional spaces characteristic of LLMs. Consequently, this approach often leads to suboptimal parameter updates, especially in scenarios where parameter interactions are significant, such as in the complex architectures of large language models (Zhang et al., 2024a).
@@
  • Another limitation of Adam and its variants lies in the substantial memory requirement for storing per-parameter adaptive learning rates and gradient statistics. As LLM sizes increase, it becomes prohibitively large, impeding scalability.
@@
To address the limitations of Adam and its variants, researchers have pursued two main approaches. The first approach, exemplified by algorithms such as Shampoo (Gupta et al., 2018) and the more recent SOAP (Vyas et al., 2024), employs sophisticated techniques to capture curvature information and parameter interdependencies. These methods utilize rotational matrices, derived through (approximate) singular value decomposition (SVD) of the gradient matrix, to provide a more comprehensive representation of the loss landscape’s geometry. This approach allows for a better approximation of the full preconditioning matrix, enabling the capture of inter-coordinate dependencies. However, the improved capability of representing parameter interactions comes at the cost of substantial computational and memory overhead (approximately twice of memory usage of Adam for transformers), rendering these algorithms challenging to implement for large-scale LLMs, where memory efficiency is crucial.
@@
The second approach focuses on reducing memory consumption through various approximation techniques. Algorithms such as AdaFactor (Shazeer and Stern, 2018) and Adam-mini (Zhang et al., 2024b) aim to decrease memory usage by approximating the second moment of the gradient matrix. Adam-mini employs a component-specific approach, averaging second moments neuron-wise for certain layers (e.g., output embeddings, value projections, and MLP layers) and head-wise for others (e.g., query and key projections). Meanwhile, AdaFactor utilizes a rank-1 approximation of the second moments. While these methods effectively reduce memory consumption, their approximations often oversimplify the complex structure of the gradient matrix’s moments, leading to a significant loss of critical curvature information and potentially compromising optimization performance. The trade-off between memory efficiency and the preservation of essential gradient statistics remains a crucial challenge in the development of scalable optimizers for LLMs.
@@
More recent approaches, such as GaLore (Zhao et al., 2024a) and MUON (Jordan et al., 2024), have attempted to strike a balance between computational complexity, memory consumption, and optimization performance in LLM training. GaLore, which can be viewed as a memory-efficient variant of SOAP, approximates the first and second moments of the gradient matrix in the leading eigensubspace. While this approach effectively reduces memory consumption, Liang et al. (2024) find that its effectiveness diminishes for sequence lengths exceeding 256. MUON, essentially an approximation of Shampoo based on some Newton-Schulz transformation proposed in Bernstein and Newhouse (2024), aims to decrease computational complexity. However, this algorithm tends to overfit to the eigensubspaces of the gradient matrix at the current iteration, failing to account for their dynamic nature throughout the optimization process. These methods, while innovative, highlight the ongoing challenge of developing optimizers that are both scalable and effective for LLMs.
@@
In this paper, we propose COSMOS, a novel hybrid optimizer that addresses the limitations of existing methods by exploiting the varying importance of eigensubspaces in the gradient matrix. Our approach decomposes the gradient into two parts: a projection onto the leading eigensubspace and a projection onto the remaining eigensubspace. The leading eigensubspace captures the most significant directions of change in the gradient, typically corresponding to the most important optimization dynamics. For this part, we apply a SOAP-like optimization strategy, tailored specifically to this reduced-dimensional space. The remaining eigensubspace, while less critical, still significantly influences optimization performance.
To address this, we employ MUON as a more efficient alternative to SOAP for this high-dimensional space. Such a hybrid approach allows COSMOS to maintain optimization effectiveness while significantly reducing memory requirements, potentially enabling the training of even larger LLMs or the use of increased batch sizes. Central to our method are (1) the exponential moving averages (EMAs) of the gradient matrix’s first and second moments, which provide smoothed estimates of recent gradient statistics, enhancing the stability of parameter updates and (2) an efficient one-step power update to identify the top $k$ eigenvectors (where $k$ is usually small, only $5\%$$10\%$ of the original dimension) that capture these crucial gradient directions, designating them as the leading eigensubspace. The design of COSMOS is motivated by empirical insights and practical considerations gleaned from extensive experimentation with large language models, offering a promising solution for scaling up LLM training.
We highlight the key contributions of this paper as follows:
@@
  • We propose a novel hybrid optimization strategy. This leads us to develop the COSMOS algorithm, which synergizes the strengths of SOAP and MUON by decomposing the gradient matrix into eigensubspaces of varying importance.
  • COSMOS achieves significant memory consumption reduction compared to the SOAP algorithm, while achieving equally or better optimization performance.
  • Unlike GaLore, COSMOS’s performance does not degrade when the sequence length exceeds 256.
The rest of the paper is organized as follows: Section 2 reviews the related work; Section 3 presents the COMOS algorithm; Section 4 presents our extensive numerical experiments; Section 5 draws a brief conclusion.
@@ 2 Related Work
Coordinate-wise Adaptive Optimizers: Adam (Kingma, 2014) and its variant AdamW (Loshchilov, 2017) have become de facto standards in deep learning optimization due to their ability to dynamically adjust learning rates based on the first and second moments of the gradients. However, these methods treat parameters independently, failing to capture interdependencies between coordinates. This limitation can lead to suboptimal updates, especially in the complex architectures of LLMs. Other adaptive optimizers such as Lion (Chen et al., 2023), Sophia (Liu et al., 2023), and Adafactor (Shazeer and Stern, 2018; Zhai et al., 2022) have shown comparable performance to AdamW in LLM pretraining but have not significantly surpassed it, suggesting the need for non-diagonal preconditioners.
Second-Order Optimizers: Researchers have explored second-order optimization techniques for training large models. These methods can be broadly categorized into Hessian-free approaches and Hessian estimation methods. Hessian-free methods, such as those proposed by Martens (2010) and Martens and Grosse (2015), optimize without explicitly computing the Hessian matrix. On the other hand, Hessian estimation methods maintain an efficient approximation of the Hessian for neural networks. Notable examples include KFAC (Martens and Grosse, 2015) and Shampoo (Gupta et al., 2018).
  • KFAC and Its Variants: KFAC (Martens and Grosse, 2015) was one of the first approaches to go beyond diagonal preconditioners in neural networks, demonstrating that a layer-wise Kronecker-factored preconditioner approximates the layer-wise Hessian in multi-layer perceptrons (MLPs). Subsequent works (Martens et al., 2018; Osawa et al., 2018) extended KFAC to other architectures. Recent research (George et al., 2018; Gao et al., 2021) has further improved trace and diagonal estimates for KFAC. Efforts to scale up KFAC (Ba et al., 2017; Puiu, 2022b, a; Eschenhagen et al., 2023) have focused on making the inversion step more efficient or enhancing distributed implementations.
  • Shampoo and Its Variants: Shampoo (Gupta et al., 2018), another second-order optimization algorithm, is motivated by the online learning algorithm Adagrad (Duchi et al., 2011). Shampoo also employs a layer-wise Kronecker-factored preconditioner. A recent distributed implementation of Shampoo (Shi et al., 2023) won an optimization efficiency benchmark (Dahl et al., 2023), highlighting the practical utility of second-order methods in deep learning. Other works (Anil et al., 2020; Peirson et al., 2022; Lin et al., 2024; Wang et al., 2024; Zhao et al., 2024b) have proposed various strategies to improve Shampoo’s scalability.
  • SOAP: The Shampoo with Adam in the Preconditioner’s eigenbasis (SOAP) algorithm (Vyas et al., 2024) establishes a formal connection between Shampoo and Adafactor. SOAP is equivalent to running Adafactor in the eigenbasis of Shampoo’s preconditioner, leading to a simpler and computationally efficient algorithm. By continually updating the running average of the second moment in the current (slowly changing) coordinate basis, SOAP mitigates the performance degradation associated with less frequent eigendecomposition computations. SOAP has shown significant improvements over AdamW in per-token efficiency.
@@
  • Adam-mini: Zhang et al. (2024b) achieve comparable performance than AdamW with a 50% smaller memory footprint. It reduces memory by carefully partitioning parameters into blocks and assigning a single learning rate to each block based on the Hessian structure of neural networks.
  • Adafactor: Shazeer and Stern (2018) use a low-rank approximation of the second moments to reduce memory consumption. It has been widely used in transformer-based models due to its memory efficiency.
  • GaLore: Zhao et al. (2024a) reduce Adam’s memory footprint by maintaining momentum in a low-rank subspace derived from the singular value decomposition (SVD) of the gradients. However, its effectiveness diminishes for sequence lengths exceeding 256, as shown in Liang et al. (2024).
  • MUON: The MUON optimizer (Jordan et al., 2024) can be viewed as an efficient approximation of Shampoo. It employs a Newton-Schulz transformation to approximately implement the Kronecker-factored preconditioner. While computationally more complex than Adam, MUON only adds minor overhead to the overall training time due to efficient parallelization of matrix operations.
These advancements in optimization techniques highlight the ongoing efforts to improve the training efficiency and performance of LLMs. However, each approach comes with its own trade-offs in terms of computational complexity, memory requirements, and optimization performance. Our work builds upon these insights to develop a hybrid approach that aims to balance these factors effectively, combining the strengths of different methods to achieve both memory efficiency and robust optimization performance for massive LLMs.
@@ 3 COSMOS: A Hybrid Adaptive Optimizer
We present a novel hybrid optimizer – COSMOS, which can achieve memory efficiency without compromising optimization performance for training LLMs. Without loss of generality, we use $m$ and $n$ to denote the numbers of rows and columns in a $m$ by $n$ matrix, and we assume $m>n$. For simplicity, we use the following notations.
$\bullet$ Matrix Sign Operator (Roberts, 1980): Given a matrix $X\in\mathbb{R}^{m\times n}$, we consider its SVD decomposition $X=UDV^{\top}$, where $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix containing all singular values of $X$, and $U\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times n}$ are left and right singular vector matrices, respectively. We define
@@
The Shampoo algorithm uses the matrix sign operator to normalize the first moment of the stochastic gradient.
@@
[…]
where $X_{5}$ is obtained by
$X_{k+1}=aX_{k}+bX_{k}X_{k}^{\top}X_{k}+cX_{k}X_{k}^{\top}X_{k}X_{k}^{\top}X_{k}$
for $k=0,1,...,4$ with $a=3.4445$, $b=-4.7750$ and $c=2.0315$. The NS transformation was mentioned in Bernstein and Newhouse (2024) to approximate the matrix sign operator without specifying the coefficient. Jordan et al. (2024) later used an ad-hoc gradient based approach to find such a set of coefficients.
$\bullet$ Normalization operator:
$\texttt{NORM}(X)=\sqrt{n}X/\|X\|_{\rm F},$ (1)
where $\|\cdot\|_{\rm F}$ denotes the Frobenius norm. The normalization operator is used to normalize the output of the NS transformation.
@@
$\bullet$ Compute the top-$r$ eigenvectors: $\texttt{SVD}(X,r)$.
We summarize COSMOS in Algorithm 1.
@@
11:     $W_{t+1}\leftarrow W_{t}-\eta\cdot\texttt{NORM}(\tilde{G}_{t})$
@@
We have two remarks regarding the differences between COSMOS and SOAP (Algorithm 2111To fairly compare with COSMOS, we only consider the one-sided SOAP, which is discussed in Section 7.1 of Vyas et al. (2024).):
Remark 1.
Lines 4–8 in Algorithm 1 can be viewed as a low-rank variant of SOAP with significantly lower memory usage. Specifically, Line 4 obtains the Exponential Moving Average (EMA) of $G_{t}$ – denoted by $M_{t}$. Line 5 approximates the leading eigenbases of $M_{t}$ – denoted by $U_{t}$ – by running a simple power iteration on the approximate EMA of $G_{t}^{\top}G_{t}$. Note that we maintain the low dimensional projection of the EMA of $G_{t}^{\top}G_{t}$ – denoted by $S_{t}$ as in Line 6. When computing the approximate EMA of $G_{t}^{\top}G_{t}$, we can project $S_{t}$ back to the original space. Line 7 then obtains the EMA of the second moment of the stochastic gradient projected onto the low dimensional eigensubspace of $M_{t}$ – denoted by $V_{t}$. Line 8 first applies an Adam-like adaptive update in the low dimensional eigensubspace of $M_{t}$, and then projects the update back to the original space – denoted by $A_{t}$.
As can be seen, COSMOS only needs to maintain four matrices in the memory: $M_{t}\in\mathbb{R}^{m\times n}$, $U_{t}\in\mathbb{R}^{n\times r}$, $S_{t}\in\mathbb{R}^{r\times r}$ and $V_{t}\in\mathbb{R}^{m\times r}$. In sharp contrast, SOAP needs to maintain $M_{t}\in\mathbb{R}^{m\times n}$, $L_{t}\in\mathbb{R}^{n\times n}$, $U_{t}\in\mathbb{R}^{n\times n}$ and $V_{t}\in\mathbb{R}^{m\times n}$. The resulting memory overhead is which is significantly larger than that of COSMOS.
Remark 2.
Line 9 in Algorithm 1 is a MUON-like update – denoted by $B_{t}$, but only applied to the complementary subspace of $M_{t}$ – denoted by $M_{t}(I-U_{t}U_{t}^{\top})$. Line 10 obtains a weighted combination of $B_{t}$ and the SOAP-like update $A_{t}$. Line 11 then applies such a combination to update the weight.
The NS transformation does not need to maintain any additional matrix in the memory, and therefore incurs no additional memory overhead.
0:  Learning rate $\eta$, momentum parameters $(\beta_{1},\beta_{2})$, perturbation parameter $\epsilon$. For simplicity, we omit the initialization.
@@
5:     $L_{t}\leftarrow\beta_{2}G_{t}^{\top}G_{t}U_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t}U_{t-1}$
@@
Algorithm 2 (One-side) SOAP for an $m\times n$ layer $W$. Per layer, we maintain four matrices: $U_{t}\in\mathbb{R}^{n\times n}$, $L\in\mathbb{R}^{n\times n}$, $V\in\mathbb{R}^{m\times n}$ and $M\in\mathbb{R}^{m\times n}$.
We next compare COSMOS with MUON (Algorithm 3) and have the following two remarks:
Remark 3.
The NS transformation in MUON behaves like a fixed point iteration without changing the eigensubspace of $M_{t}$. Therefore, the resulting preconditioning is dedicated to the eigenbases of $M_{t}$. In contrast, COSMOS adopts a power iteration to the approximate EMA of $G_{t}^{\top}G_{t}$ for incrementally updating the projection matrix $U_{t}$ from $U_{t-1}$. Therefore, we can separately control the training dynamics of $U_{t}$, yielding a smoother transition of eigenbases for preconditioning, which potentially leads to better optimization performance.
Remark 4.
We propose a new normalization procedure NORM in (1). It is motivated by the observation that NS5 is only an approximation of the standard Newton Schulz iteration, and the singular values of its output are not exactly $1$’s, especially for small singular values. Therefore, the Frobenius norm of the output may substantially deviate from the desirable $\sqrt{n}$ (corresponding to the case where all singular values equal to $1$). In contrast, our normalization procedure guarantees the Frobenius norm of the output to be exactly $\sqrt{n}$, which improves training stability.
0:  Learning rate $\eta$, momentum parameters $\beta_{1}$. For simplicity, we omit the initialization.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable success across various domains, yet their optimization remains a significant challenge due to the complex and high-dimensional loss landscapes they inhabit. While adaptive optimizers such as AdamW are widely used, they suffer from critical limitations, including an inability to capture interdependencies between coordinates and high memory consumption. Subsequent research, exemplified by SOAP, attempts to better capture coordinate interdependence but incurs greater memory overhead, limiting scalability for massive LLMs. An alternative approach aims to reduce memory consumption through low-dimensional projection, but these methods lose the gradient information in the residual space, resulting in less effective optimization. In this paper, we propose COSMOS, a novel hybrid optimizer that leverages the varying importance of eigensubspaces in the gradient matrix to achieve memory efficiency without compromising optimization performance. The design of COSMOS is motivated by our empirical insights and practical considerations. Specifically, COSMOS applies SOAP to the leading eigensubspace, which captures the primary optimization dynamics, and MUON to the remaining eigensubspace, which is less critical but computationally expensive to handle with SOAP. This hybrid strategy significantly reduces memory consumption while maintaining robust optimization performance, making it particularly suitable for massive LLMs. Numerical experiments on various datasets and transformer architectures are provided to demonstrate the effectiveness of COSMOS. Our code is available at https://github.com/lliu606/COSMOS.
@@ 1 Introduction
The optimization of Large Language Models (LLMs) is fundamental to their success, enabling these models to achieve state-of-the-art performance across diverse tasks. However, the non-convex loss landscapes inherent to LLMs, which can contain hundreds of billions or even trillions of parameters (Achiam et al., 2023), present significant optimization challenges. Adaptive optimizers, such as Adam (Kingma, 2014) and its variants AdamW (Loshchilov, 2017), have emerged as de facto standards due to their ability to dynamically adjust learning rates based on the second moment of the gradient. Despite their widespread adoption, these methods suffer from two critical limitations that impede their effectiveness and scalability in the context of increasingly large and complex LLMs:
(I) Adam and its variants have limitations in adaptive learning rates. By adjusting learning rates independently for each parameter, the method reduces computational complexity but may not fully capture parameter interdependencies. In complex architectures of LLMs, this independent approach can lead to suboptimal parameter updates (Zhang et al., 2024a).
@@
(II) Another limitation of Adam and its variants lies in the substantial memory requirement for storing per-parameter adaptive learning rates and gradient statistics. As LLM sizes increase, memory consumption becomes prohibitively large, impeding scalability.
@@
To address the limitations of Adam and its variants, researchers have pursued two main approaches. The first approach, exemplified by algorithms such as Shampoo (Gupta et al., 2018) and the more recent SOAP (Vyas et al., 2024), employs sophisticated techniques to capture curvature information and parameter interdependencies. These methods utilize rotational matrices, derived through (approximate) singular value decomposition (SVD) of the gradient matrix, to provide a more comprehensive representation of the loss landscape’s geometry. This approach allows for a better approximation of the full preconditioning matrix, enabling the capture of inter-coordinate dependencies. However, the improved capability of representing parameter interactions comes at the cost of substantial computational and memory overhead, rendering these algorithms challenging to implement for large-scale LLMs, where memory efficiency is crucial.
@@
The second approach focuses on reducing memory consumption through various approximation techniques. Algorithms such as AdaFactor (Shazeer and Stern, 2018) and Adam-mini (Zhang et al., 2024b) aim to decrease memory usage by approximating the second moment of the gradient matrix. Adam-mini employs a component-specific approach, averaging second moments neuron-wise for certain layers. Meanwhile, AdaFactor utilizes a rank-1 approximation of the second moments. While these methods reduce memory cost, their approximations oversimplify the structure of the gradient matrix’s second order moments, compromising optimization performance. The trade-off between memory efficiency and the preservation of gradient statistics remains a crucial challenge.
@@
More recent approaches, such as GaLore (Zhao et al., 2024a) and MUON (Jordan et al., 2024), have attempted to strike a balance between computational complexity, memory consumption, and optimization performance in LLM training. GaLore, which can be viewed as a memory-efficient variant of SOAP, approximates the first and second moments of the gradient matrix in the leading eigensubspace. While it effectively reduces memory consumption, Liang et al. (2024) find that its effectiveness diminishes for sequence lengths exceeding 256. MUON, essentially an approximation of Shampoo based on Newton-Schulz transformation proposed in Bernstein and Newhouse (2024), aims to decrease computational complexity. However, MUON only estimates the eigensubspaces based on the gradient on one batch, rather than capturing the comprehensive distribution of gradients across the entire optimization process.
@@
In this paper, we propose COSMOS, a novel hybrid optimizer that addresses the limitations of existing methods by exploiting the varying importance of eigensubspaces in the gradient matrix. Our approach decomposes the gradient into two parts: a projection onto the leading eigensubspace and a projection onto the remaining eigensubspace. The leading eigensubspace captures the most significant directions of change in the gradient, typically corresponding to the most important optimization dynamics. For this part, we apply a SOAP-like optimization strategy. However, by crucially restricting SOAP to the leading eigensubspace, COSMOS only needs to maintain the projection matrix and the second-order moment within this small subspace, thereby retaining SOAP’s ability to capture parameter interdependencies while substantially lowering its memory cost. The remaining eigensubspace, while less critical, still significantly influences optimization performance. To address this, we employ MUON as a more efficient alternative to SOAP for this high-dimensional space. Such a hybrid approach allows COSMOS to maintain optimization effectiveness while significantly reducing memory requirements compared to SOAP, potentially enabling the training of larger LLMs or the use of increased batch sizes.
@@
We highlight the key contributions of this paper as follows: (1) We propose a novel hybrid optimization strategy. This leads us to develop the COSMOS algorithm, which synergizes the strengths of SOAP and MUON by decomposing the gradient matrix into eigensubspaces of varying importance. (2) COSMOS achieves significant memory consumption reduction compared to the SOAP algorithm, while achieving equally or better optimization performance.
@@ 2 Related Work
Coordinate-wise adaptive optimizers: Adam (Kingma, 2014) and AdamW (Loshchilov, 2017) have become standards in deep learning optimization due to their ability to dynamically adjust learning rates based on the first and second moments of the gradients. However, these methods treat parameters independently, failing to capture interdependencies between coordinates. This limitation can lead to suboptimal updates, especially in the complex architectures of LLMs. Other adaptive optimizers such as Lion (Chen et al., 2023), Sophia (Liu et al., 2023), and Adafactor (Shazeer and Stern, 2018; Zhai et al., 2022) have shown comparable performance to AdamW in LLM pretraining but have not significantly surpassed it, suggesting the need for non-diagonal preconditioners.
Second-Order Optimizers: Researchers have explored second-order optimization techniques for training large models. These methods can be broadly categorized into Hessian-free approaches and Hessian estimation methods. Hessian-free methods, such as those proposed by Martens (2010) and Martens and Grosse (2015), optimize without explicitly computing the Hessian matrix. On the other hand, Hessian estimation methods maintain an efficient approximation of the Hessian for neural networks. Notable examples include KFAC (Martens and Grosse, 2015), Shampoo (Gupta et al., 2018) and SOAP (Vyas et al., 2024).
$\diamond$ Shampoo and Its Variants: Shampoo (Gupta et al., 2018), another second-order optimization algorithm, is motivated by the online learning algorithm Adagrad (Duchi et al., 2011). Shampoo also employs a layer-wise Kronecker-factored preconditioner. A recent distributed implementation of Shampoo (Shi et al., 2023) won an optimization efficiency benchmark (Dahl et al., 2023), highlighting the practical utility of second-order methods in deep learning. Other works (Anil et al., 2020; Peirson et al., 2022; Lin et al., 2024; Wang et al., 2024; Zhao et al., 2024b) have proposed various strategies to improve Shampoo’s scalability.
$\diamond$ SOAP: SOAP algorithm (Vyas et al., 2024) establishes a formal connection between Shampoo and Adafactor. SOAP is equivalent to running Adafactor in the eigenbasis of Shampoo’s preconditioner, leading to a simpler and computationally efficient algorithm. By continually updating the running average of the second moment in the current (slowly changing) coordinate basis, SOAP mitigates the performance degradation associated with less frequent eigendecomposition computations. SOAP has shown significant improvements over AdamW in per-token efficiency.
@@
$\diamond$ Adafactor and Adam-mini: Shazeer and Stern (2018) use a low-rank approximation of the second moments to reduce memory consumption. It has been widely used in LLMs due to memory efficiency. Zhang et al. (2024b) achieve comparable performance than AdamW with a 50% smaller memory. It reduces memory by carefully partitioning parameters into blocks and assigning a single learning rate to each block based on the Hessian structure of neural networks.
$\diamond$ GaLore: Zhao et al. (2024a) reduce Adam’s memory by maintaining momentum in a low-rank subspace derived from the singular value decomposition (SVD) of the gradients. However, its effectiveness diminishes for sequence lengths exceeding 256, as shown in Liang et al. (2024).
$\diamond$ MUON: MUON optimizer (Jordan et al., 2024) can be viewed as an efficient approximation of Shampoo. It employs Newton-Schulz transformation to approximately implement the Kronecker-factored preconditioner. While computationally more complex than Adam, it only adds minor overhead to the overall training time due to efficient parallelization of matrix operations.
These advancements highlight the efforts to improve training efficiency and performance of LLMs. However, each approach comes with its own trade-offs in terms of computational complexity, memory requirements, and performance. Our work builds upon these insights to develop a hybrid approach that aims to balance these factors effectively, combining the strengths of different methods to achieve both memory efficiency and better optimization performance for LLMs.
@@ 3 COSMOS: A Hybrid Adaptive Optimizer
We present a novel hybrid optimizer – COSMOS in Algorithm 1, which can achieve memory efficiency without compromising performance for training LLMs. Without loss of generality, we use $m$ and $n$ to denote the numbers of rows and columns in a $m$ by $n$ matrix, and we assume $m>n$. For simplicity, we use the following notations:
$\bullet$ Matrix Sign Operator: Given a matrix $X\in\mathbb{R}^{m\times n}$ and its reduced-SVD $X=UDV^{\top}$, where $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix containing all singular values of $X$, and $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{n\times n}$ are left and right singular vector matrices, respectively. We define
@@
$\texttt{NS5}(X_{0})=X_{5}.$
where $X_{5}$ is obtained by $X_{k+1}=aX_{k}+bX_{k}X_{k}^{\top}X_{k}+cX_{k}X_{k}^{\top}X_{k}X_{k}^{\top}X_{k}$ for k=0,1,,4k=0,1,...,4 with $a=3.4445$, b=4.7750b=-4.7750 and $c=2.0315$. Bernstein and Newhouse (2024) first mentioned this transformation to approximate the matrix sign operator without specifying the coefficient. Jordan et al. (2024) later used an ad-hoc gradient based approach to find the set of coefficients here.
$\bullet$ Normalization operator: $\texttt{NORM}(X)=\sqrt{n}X/\|X\|_{\rm F}$, where $\|\cdot\|_{\rm F}$ denotes the Frobenius norm. The normalization operator is used to normalize the output of the NS transformation.
@@
11:  $W_{t+1}\leftarrow W_{t}-\eta\cdot\texttt{NORM}(\tilde{G}_{t})\cdot\sqrt{m}$
@@
Design Principle: The design of COSMOS is guided by a simple principle: instead of maintaining SOAP’s full second-moment matrix—which is memory-prohibitive—we track its dominant eigenspace and operate in a low-dimensional subspace.
In the SOAP algorithm, the exponential moving average (EMA) of the second moment is
$H_{t}=\beta_{2}H_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t},$ (1)
where $G_{t}$ is the stochastic gradient. Because $H_{t}\in\mathbb{R}^{n\times n}$ is dense, storing it is infeasible for large $n$. COSMOS avoids this by maintaining (i) an orthonormal basis $U_{t}\in\mathbb{R}^{n\times r}$ for the leading eigenspace of $H_{t}$ and (ii) a projected second-moment matrix $S_{t}\in\mathbb{R}^{r\times r}$ with
$S_{t}\approx U_{t}^{\top}H_{t}U_{t}.$
Assume at step t1t-1 that $U_{t-1}$ spans the dominant eigenspace and $S_{t-1}\approx U_{t-1}^{\top}H_{t-1}U_{t-1}$. Approximating $H_{t-1}$ by its rank-$r$ surrogate, $U_{t-1}S_{t-1}U_{t-1}^{\top}$, and substituting into Equation 1 yields
$\tilde{H}_{t}=\beta_{2}\,U_{t-1}S_{t-1}U_{t-1}^{\top}+(1-\beta_{2})\,G_{t}^{\top}G_{t}.$
We then update the basis via a one-step power iteration:
$U_{t}=\texttt{QR}\big(\tilde{H}_{t}\,U_{t-1}\big),$
and refresh the projected second moment by
$S_{t}=U_{t}^{\top}\tilde{H}_{t}U_{t}.$
These two steps track the dominant eigenspace and its curvature information with $O(nr)$ memory.
Given $U_{t}$, Line 7 of Algorithm 1 maintains the EMA of the projected gradients $V_{t}\in\mathbb{R}^{m\times r}$, and Line 8 performs a SOAP-like adaptive update within the subspace spanned by $U_{t}$, producing $A_{t}$ after projecting back to the full parameter space. This is the SOAP component of COSMOS.
To complement the low-rank update, COSMOS applies a MUON-inspired preconditioner on the orthogonal complement of $U_{t}$. Writing the orthogonal projector as $P_{t}^{\perp}=I-U_{t}U_{t}^{\top}$, Line 9 forms
$B_{t}=\texttt{NORM}\!\left(\texttt{NS5}\!\left(\frac{M_{t}P_{t}^{\perp}}{\|M_{t}P_{t}^{\perp}\|_{\mathrm{F}}}\right)\right)=\texttt{NORM}\!\left(\texttt{NS5}\!\left(\frac{M_{t}-M_{t}U_{t}U_{t}^{\top}}{\|M_{t}-M_{t}U_{t}U_{t}^{\top}\|_{\mathrm{F}}}\right)\right),$ (2)
where NS5 applies directly to the residual momentum; no additional matrices are stored.
Finally, Lines 10–11 combine the two components:
$\tilde{G}_{t}=A_{t}+\gamma B_{t}\sqrt{m},\qquad W_{t+1}=W_{t}-\eta\,\texttt{NORM}(\tilde{G}_{t})\,\sqrt{m}.$
[…]
0: Learning rate $\eta$, momentum parameters $(\beta_{1},\beta_{2})$, perturbation parameter $\epsilon$.
@@
5:  $L_{t}\leftarrow\beta_{2}G_{t}^{\top}G_{t}+(1-\beta_{2})G_{t}^{\top}G_{t}$
@@
Algorithm 2 (One-side) SOAP
0: Learning rate $\eta$, momentum parameters $\mu$.